# publications.bib

### v3

@@ 3. Mantovan’s Formula and the Hodge-Tate Period Morphism
For $K\subset\mathbf{G}(\mathbb{A}_{f})$ a sufficiently small open compact, we define $\mathcal{S}_{K}:=(\mathrm{Sh}(\mathbf{G},X)_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ to be the adic space over $\mathop{\rm Spa}(E_{\mathfrak{p}})$ attached to the Shimura variety. When $K=K_{p}^{\mathrm{hs}}K^{p}$ with $K_{p}^{\mathrm{hs}}$ a hyperspecial subgroup, the space $\mathcal{S}_{K}$ has a canonical integral model $\mathfrak{S}_{K}$ over $\mathcal{O}_{E,\mathfrak{p}}$. Let $\mathcal{S}_{K}^{\circ}\subset\mathcal{S}_{K}$ be the good reduction locus, i.e. the open subspace of $\mathcal{S}_{K}$ obtained from the adic generic fiber of the $p$-adic completion $\mathfrak{S}^{\wedge}_{K}$ of the scheme $\mathfrak{S}_{K}$. We define $\mathcal{S}_{K^{\prime}}^{\circ}\subset\mathcal{S}_{K^{\prime}}$ for $K^{\prime}\subset K$ by taking the preimage under the natural map from $\mathcal{S}_{K^{\prime}}$ to $\mathcal{S}_{K}$. We also consider the adic spaces $\mathcal{S}^{*}_{K}$ and $\mathcal{S}^{\mathrm{tor}}_{K}$ attatched to the minimal and toroidal compactification of the Shimura variety $\mathcal{S}^{*}_{K}$.
@@
Associated to the $G(\mathbb{R})$-conjugacy class $X$, we have a minuscule cocharacter $\mu$ of $G_{C}$ which is defined over $E_{\mathfrak{p}}$. Let $\mathcal{F}\ell_{G,\mu^{-1}}$ be the flag variety over $\mathop{\rm Spa}(C)$ associated to $\mu^{-1}$ the dominant inverse of $\mu$. Since $\mu$ is minuscule, via the Bialynicki-Birula isomorphism, when viewed as a diamond the flag variety $\mathcal{F}\ell_{G,\mu^{-1}}$ represents the following functor on $\mathrm{Perf}_{C}$. Given any $S\in\mathrm{Perf}_{C}$, $\mathcal{F}\ell_{G,\mu^{-1}}(S)$ is the set of modifications of vector bundles $\mathcal{E}\dashrightarrow\mathcal{E}_{0}$ of meromorphy $\mu$ on $X_{S}$, the relative Fargues-Fontaine curve over $S$, such that the modification occurs over the untilt of $S$ corresponding to the map $S\rightarrow\mathop{\rm Spd}(C)$.
@@
$\mathcal{S}_{K^{p}}^{\circ}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\leftarrow\\ K_{p}\end{subarray}}\mathcal{S}_{K^{p}K_{p}}^{\circ}\subset\mathcal{S}_{K^{p}}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\leftarrow\\ K_{p}\end{subarray}}\mathcal{S}_{K^{p}K_{p}}\subset\mathcal{S}_{K^{p}}^{\mathrm{tor}}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\leftarrow\\ K_{p}\end{subarray}}\mathcal{S}^{\mathrm{tor}}_{K^{p}K_{p}}\rightarrow\mathcal{S}_{K^{p}}^{*}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\leftarrow\\ K_{p}\end{subarray}}\mathcal{S}^{*}_{K^{p}K_{p}}$
be the associated perfectoid Shimura varieties. We also consider $\overline{\mathcal{S}}^{\circ}_{K^{p},C}$, the canonical compactification of the good reduction locus. This will be a subspace of $\mathcal{S}_{K^{p},C}$, since $\mathcal{S}_{K^{p},C}$ is partially proper. Caraiani-Scholze [CS1, §2.1] consider the Hodge-Tate period morphism on $\mathcal{S}_{K^{p}}$
$\pi_{\mathrm{HT}}:\mathcal{S}_{K^{p},C}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}},$
which records the relative position of the Hodge-Tate filtration associated with the $p$-divisible group. This extends [CS2, §4.1] to a Hodge-Tate period morphism on the minimal compactification
$\pi_{\mathrm{HT}}^{*}:\mathcal{S}_{K^{p},C}^{*}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}$
@@
$\pi_{\mathrm{HT}}^{\mathrm{tor}}:\mathcal{S}_{K^{p},C}^{\mathrm{tor}}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}.$
We write $\pi_{\mathrm{HT}}^{\circ}$ for the restriction to the good reduction locus, and $\overline{\pi}_{\mathrm{HT}}^{\circ}$ for the canonical compactification of $\pi_{\mathrm{HT}}^{\circ}$, where we note that this again maps to $\mathcal{F}\ell_{G,\mu^{-1}}$ as this is proper over $\mathop{\rm Spa}(C)$.
@@
    $\pi_{\mathrm{HT}}^{*}$ and $\pi_{\mathrm{HT}}^{\mathrm{tor}}$ are partially proper and qcqs; hence, proper.
@@
    $\pi_{\mathrm{HT}}$ and $\overline{\pi}_{\mathrm{HT}}^{\circ}$ are partially proper, but not always qcqs.
@@
    $\pi_{\mathrm{HT}}^{\circ}$ is qcqs, but not partially proper.
With these properties in mind, let us study the fibers of these maps. For our purposes, we will focus on the compactly supported cohomology of $\mathcal{S}_{K^{p},C}$ and in turn the sheaf $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ on $\mathcal{F}\ell_{G,\mu^{-1}}$.
Remark 3.1.
Our goal is to describe the stalks of $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ at a geometric point $x:\mathop{\rm Spa}(C,C^{+})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}$. We assume the geometric point $x$ factors through the adic Newton strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b}$ for $b\in B(G,\mu)$, and choose a completely slope divisible $p$-divisible group $\mathbb{X}_{b}$ over $\overline{\mathbb{F}}_{p}$ corresponding to $b\in B(G,\mu)$. Let $\mathrm{Ig}^{b}$ be the associated perfect Igusa variety as defined in §2.2, with toroidal compactification $\mathrm{Ig}^{b,\mathrm{tor}}$ and minimal compactification $\mathrm{Ig}^{b,*}$. Recall that we have associated perfectoid Igusa varieties, $\mathfrak{Ig}^{b}_{C},\mathfrak{Ig}^{b,\mathrm{tor}}_{C},\mathfrak{Ig}^{b,*}_{C}$, which should model the fibers of $\overline{\pi}_{\mathrm{HT}}^{\circ}$, $\pi_{\mathrm{HT}}^{\mathrm{tor}}$, and $\pi^{*}_{\mathrm{HT}}$, respectively. We let $\partial\mathfrak{Ig}_{C}^{b,*}$ and $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ be the Zariski closed subspaces attached to the boundaries $\partial\mathrm{Ig}^{b,*}$ and $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, respectively.
@@
Our goal is to show that this computes the fibers of $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ for geometric points in $\mathcal{F}\ell_{G,\mu^{-1}}^{b}$.
@@
[CS2, San] There exists a diagram of spaces of the form
@@
where the maps $i$, $\phantom{}{}^{*}i$, and $\phantom{}{}^{\mathrm{tor}}i$ are open immersions whose image contains all rank $1$ points. Moreover, the fibers $(\pi_{\mathrm{HT}}^{*})^{-1}(x)$ and $(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)$ are partially proper, so in particular $\phantom{}{}^{*}i$ and $\phantom{}{}^{\mathrm{tor}}i$ are canonical compactifications in the sense of [Ecod, Proposition 18.6].
@@
This theorem in the case where $(\mathbf{G},X)$ is of PEL type $A$ attached to a globally quasi-split unitary group of even dimension is [CS2, Theorems 2.7.2,Theorem 4.5.1], and the general case of PEL type $A$ or $C$ is proven in [San, Theorems 4.3.10,4.3.12]. ∎
@@
[CS2, San] The partially minimally compactified Igusa variety $\mathrm{Ig}^{b,*}$ is affine; in particular, the attatched adic space $\mathfrak{Ig}^{b,*}_{C}$ is affinoid perfectoid. Moreover, there exists a proper map
@@
For the affineness, the case of PEL type $A$ attached to a global quasi-split unitary group of even dimension is covered by [CS2, Theorem 1.7] and [CS2, Lemma 4.5.2]. The general case of PEL type $A$ or $C$ is covered in [San, Lemma 3.3.7]. To see the properness, we note that the map
@@
is the one appearing in the Stein factorization described in [CS2, Proposition 3.3.4] and [San, Proposition 3.3.5]. This also shows that one has an isomorphism on global sections. ∎
@@
The fact that $\mathfrak{Ig}_{C}^{b,*}$ is affinoid perfectoid follows from the previous Theorem. Moreover, $\partial\mathfrak{Ig}_{C}^{b,*}$ is a Zariski closed subspace, since it came from considering the adic generic fiber of a formal model of the perfect closed subscheme $\partial\mathrm{Ig}^{b,*}\subset\mathrm{Ig}^{b,*}$. The claim for the toroidal compactification follows since the map $\mathfrak{Ig}_{C}^{b,\mathrm{tor}}\rightarrow\mathfrak{Ig}_{C}^{b,*}$ is proper, and maps the boundary $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ to $\partial\mathfrak{Ig}_{C}^{b,*}$. ∎
@@
For $x:\mathop{\rm Spa}(C,C^{+})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{b}$ a geometric point, we have isomorphisms
@@
induced by the natural open immersions $\pi_{\mathrm{HT}}^{-1}(x)\hookrightarrow(\pi_{\mathrm{HT}}^{*})^{-1}(x)\simeq\overline{\mathfrak{Ig}}_{C}^{b,*}$ (resp. $\pi_{\mathrm{HT}}^{-1}(x)\hookrightarrow(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)\simeq\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$), as given by Theorem 3.2. Here $\overline{\partial\mathfrak{Ig}}_{C}^{b,*}$ (resp. $\overline{\partial\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$) is the Zarisiki closed subset of $\overline{\mathfrak{Ig}}_{C}^{b,*}$ (resp. $\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$) defined by the canonical compactification of the boundaries $\partial\mathfrak{Ig}^{b}_{C}\subset\mathfrak{Ig}^{b}_{C}$ (resp. $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}\subset\mathfrak{Ig}^{b,\mathrm{tor}}_{C}$).
@@
obtained by base-changing the closed immersion $\partial\mathcal{S}^{\mathrm{tor}}_{K^{p},C}\hookrightarrow\mathcal{S}^{\mathrm{tor}}_{K^{p},C}$ to the fiber $(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)$ and applying Theorem 3.2. To show the desired claim, it suffices to show this is an isomorphism. Note that both the LHS and RHS are partially proper; therefore, to show this is an isomorphism, it suffices to show it induces an isomorphism on rank $1$ points. In particular, given $C^{\prime}/C$ a complete algebraically closed non-archimedean field, we claim that there exists a Cartesian diagram of the form

### v5

@@ 3. Mantovan’s Formula and the Hodge-Tate Period Morphism
We now review some basic facts about the geometry of the Hodge-Tate period morphism, establishing some facts about the fiber of the Hodge-Tate period morphism in §3.1. In §3.2, we then apply these basic geometric facts to establish a version of Mantovan’s product formula for non-compact Shimura varieties (Theorem 5.1). This will in particular allow us to bound the degrees of the torsion cohomology of the Shimura variety in terms of the semiperversity of certain sheaves coming from Igusa varieties and the perverse t-exactness of Hecke operators on $\mathrm{Bun}_{G}$.
We recall that $(\mathbf{G},X)$ denotes a Shimura datum with reflex field $E/\mathbb{Q}$. We let $\mathrm{Sh}(\mathbf{G},X)_{K}/\mathop{\rm Spec}(E)$ denote the associated Shimura variety for a sufficiently small open compact subgroup $K\subset\mathbf{G}(\mathbb{A}_{f})$. We will assume throughout that the Shimura datum $(\mathbf{G},X)$ is of PEL type AC.
We recall that $\ell\neq p$ are distinct primes, $E_{\mathfrak{p}}$ is the completion at the place $\mathfrak{p}|p$ determined by a fixed choice of isomorphism $j:\overline{\mathbb{Q}}_{p}\simeq\mathbb{C}$, and $C$ is the completed algebraic closure of $E_{\mathfrak{p}}$. The group $G:=\mathbf{G}_{\mathbb{Q}_{p}}$ will denote the local group, and we let $K=K^{p}K_{p}$ be the decomposition into the level $K^{p}\subset\mathbf{G}(\mathbb{A}_{f}^{p})$ away from $p$ (resp. the level $K_{p}\subset G(\mathbb{Q}_{p})$ at $p$).
3.1. The Hodge-Tate Period Morphism
@@
For $K\subset\mathbf{G}(\mathbb{A}_{f})$ a sufficiently small open compact, we define $\mathcal{S}(\mathbf{G},X)_{K}:=(\mathrm{Sh}(\mathbf{G},X)_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ to be the adic space over $\mathop{\rm Spa}(E_{\mathfrak{p}})$ attached to the Shimura variety. For brevity, in this section we will often denote this simply by $\mathcal{S}_{K}$. When $K=K_{p}^{\mathrm{hs}}K^{p}$ with $K_{p}^{\mathrm{hs}}$ a hyperspecial subgroup, the space $\mathcal{S}_{K}$ has a canonical integral model $\mathscr{S}_{K}$ over $\mathcal{O}_{E,\mathfrak{p}}$. Let $\mathcal{S}_{K}^{\circ}\subset\mathcal{S}_{K}$ be the good reduction locus, i.e. the open subspace of $\mathcal{S}_{K}$ obtained from the adic generic fiber of the $p$-adic completion $\mathscr{S}^{\wedge}_{K}$ of the scheme $\mathscr{S}_{K}$. We define $\mathcal{S}_{K^{\prime}}^{\circ}\subset\mathcal{S}_{K^{\prime}}$ for $K^{\prime}\subset K$ by taking the preimage under the natural map from $\mathcal{S}_{K^{\prime}}$ to $\mathcal{S}_{K}$. We also consider the adic spaces $\mathcal{S}^{*}_{K}:=(\mathrm{Sh}(\mathbf{G},X)^{*}_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ and $\mathcal{S}^{\mathrm{tor}}_{K}:=(\mathrm{Sh}(\mathbf{G},X)^{\mathrm{tor}}_{K}\otimes_{E}E_{\mathfrak{p}})^{\mathrm{ad}}$ attached to the minimal and toroidal compactifications of the Shimura variety $\mathrm{Sh}(\mathbf{G},X)_{K}$ over $E$, respectively, as described in §2.1.3, where, for the former, we fix a compatible choice $\Sigma=\{\Sigma_{\Phi}\}$ of admissible smooth rational polyhedral cone decomposition data, as described in loc.cit. We use the superscript $(-)^{\Diamond}$ for the diamond attached to these analytic adic spaces, as defined in [55, Lecture X].
@@
Associated to the $\mathbf{G}(\mathbb{R})$-conjugacy class $X$, we have a conjugacy class of a minuscule cocharacter $\mu$ of $G_{C}$ with reflex field $E_{\mathfrak{p}}$ attached to the inverse of the Hodge cocharacter of $X$. Let $\mathcal{F}\ell_{G,\mu^{-1}}$ be the adic flag variety over $\mathop{\rm Spa}(C)$ associated to the conjugacy class of $\mu^{-1}$, the dominant inverse of $\mu$ (i.e we quotient out by the parabolic defined by the set of $g\in G_{C}$ such that $\mathop{\rm lim}\nolimits_{t\rightarrow 0}\mathrm{ad}(\mu(t))g$ exists). Since $\mu$ is minuscule, via the Bialynicki-Birula isomorphism ([7, Theorem 3.4.5]), the diamond attached to the flag variety $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ represents the following functor on $\mathrm{Perf}_{C}$. Given any $S\in\mathrm{Perf}_{C}$, $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}(S)$ is the set of modifications of vector bundles $\mathcal{E}\dashrightarrow\mathcal{E}_{0}$ of meromorphy $\mu$ on $X_{S}$, the relative Fargues-Fontaine curve over $S$, such that the modification occurs over the degree one Cartier divisor in $X_{S}$ corresponding to the untilt of $S$ defined by the map $S\rightarrow\mathop{\rm Spd}(C)$.
@@
$\mathcal{S}_{K^{p},C}^{\circ}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ K_{p}\end{subarray}}\mathcal{S}_{K^{p}K_{p},C}^{\circ,\Diamond}\subset\mathcal{S}_{K^{p},C}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ K_{p}\end{subarray}}\mathcal{S}_{K^{p}K_{p},C}^{\Diamond}\subset\mathcal{S}_{K^{p},C}^{\mathrm{tor}}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ K_{p}\end{subarray}}\mathcal{S}^{\mathrm{tor},\Diamond}_{K^{p}K_{p},C}\rightarrow\mathcal{S}_{K^{p},C}^{*}:=\mathop{\rm lim}\nolimits\limits_{\begin{subarray}{c}\longleftarrow\\ K_{p}\end{subarray}}\mathcal{S}^{*,\Diamond}_{K^{p}K_{p},C}$
be the associated perfectoid Shimura varieties [^4] If $(\mathbf{G},X)$ is more generally of Hodge type then this is representable by a perfectoid space by [57, Theorem IV.1.1], and in general they have the structure of diamonds (See [28] and [50, Section 2]), where this limit is computed in the category of sheaves on $\mathrm{Perf}_{C}$ with respect to the $v$-topology on $\mathrm{Perf}_{C}$, as defined in [56, Definition 8.1] and we recall that any diamond defines such a $v$-sheaf in light of [56, Proposition 11.9]. Throughout, we identify adic spaces attached to perfectoids and their associated $v$-sheaves under the Yoneda embedding, recalling that the $v$-topology is sub-canonical, by [56, Theorem 1.2].
We also consider $\overline{\mathcal{S}}^{\circ}_{K^{p},C}$, the canonical compactification of the good reduction locus over $\mathop{\rm Spd}(C)$ in the sense of [56, Proposition 18.6]. This will be a subspace of $\mathcal{S}_{K^{p},C}$, since $\mathcal{S}_{K^{p},C}$ is partially proper. Caraiani-Scholze [7, §2.1] consider the Hodge-Tate period morphism on $\mathcal{S}_{K^{p}}$
$\pi_{\mathrm{HT}}:\mathcal{S}_{K^{p},C}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond},$
which records the relative position of the Hodge-Tate filtration associated with the abelian varieties that $\mathcal{S}_{K^{p},C}$ parametrizes. This extends [8, §4.1] to a Hodge-Tate period morphism on the minimal compactification
$\pi_{\mathrm{HT}}^{*}:\mathcal{S}_{K^{p},C}^{*}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$
@@
$\pi_{\mathrm{HT}}^{\mathrm{tor}}:\mathcal{S}_{K^{p},C}^{\mathrm{tor}}\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}.$
We write $\pi_{\mathrm{HT}}^{\circ}$ for the restriction to the good reduction locus, and $\overline{\pi}_{\mathrm{HT}}^{\circ}$ for the canonical compactification of $\pi_{\mathrm{HT}}^{\circ}$, where we note that this again maps to $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$ as this is proper over $\mathop{\rm Spa}(C)$.
@@
    $\overline{\pi}_{\mathrm{HT}}^{\circ}$, $\pi_{\mathrm{HT}}^{*}$, and $\pi_{\mathrm{HT}}^{\mathrm{tor}}$ are partially proper and qcqs; hence, proper.
@@
    $\pi_{\mathrm{HT}}$ is partially proper, but not always qcqs (Unless the Shimura variety is compact).
@@
    $\pi_{\mathrm{HT}}^{\circ}$ is qcqs, but not partially proper (Unless the Shimura variety is compact).
With these properties in mind, let us study the fibers of these maps. For our purposes, we will focus on the compactly supported cohomology of $\mathcal{S}_{K^{p},C}$ and in turn the sheaf $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ on $\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$.
Our goal is to describe the stalks of $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ at a geometric rank $1$ point $x:\mathop{\rm Spa}(C,O_{C})\rightarrow\mathcal{F}\ell^{\Diamond}_{G,\mu^{-1}}$. We assume the geometric point $x$ factors through the adic Newton strata $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ for $b\in B(G,\mu)$, and choose a completely slope divisible $p$-divisible group $\mathbb{X}_{b}$ over $\overline{\mathbb{F}}_{p}$ corresponding to $b\in B(G,\mu)$. Let $\mathrm{Ig}^{b}$ be the associated perfect Igusa variety as defined in §2.2.1, with toroidal compactification $\mathrm{Ig}^{b,\mathrm{tor}}$ and minimal compactification $\mathrm{Ig}^{b,*}$. Recall that we have associated perfectoid Igusa varieties, $\mathfrak{Ig}^{b}_{C}$$\mathfrak{Ig}^{b,\mathrm{tor}}_{C}$, and $\mathfrak{Ig}^{b,*}_{C},$ which should model the fibers of $\overline{\pi}_{\mathrm{HT}}^{\circ}$, $\pi_{\mathrm{HT}}^{\mathrm{tor}}$, and $\pi^{*}_{\mathrm{HT}}$, respectively. We let $\partial\mathfrak{Ig}_{C}^{b,*}$ and $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ be the Zariski closed subspaces attached to the boundaries $\partial\mathrm{Ig}^{b,*}$ and $\partial\mathrm{Ig}^{b,\mathrm{tor}}$, respectively.
@@
Our goal is to show that this computes the fibers of $R\pi_{\mathrm{HT}!}(\overline{\mathbb{F}}_{\ell})$ at geometric points of $\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$.
@@
[8, 54] For a geometric rank $1$ point, $x:\mathop{\rm Spa}(C,O_{C})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{\Diamond}$, there exists a diagram of spaces of the form
@@
where the maps $i$, $\phantom{}{}^{*}i$, and $\phantom{}{}^{\mathrm{tor}}i$ are open immersions whose image contains all rank $1$ points. Moreover, the fibers $(\pi_{\mathrm{HT}}^{*})^{-1}(x)$ and $(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)$ are partially proper, so in particular $\phantom{}{}^{*}i$ and $\phantom{}{}^{\mathrm{tor}}i$ are canonical compactifications in the sense of [56, Proposition 18.6].
@@
This theorem in the case where $(\mathbf{G},X)$ is of PEL type $A$ attached to a globally quasi-split unitary group of even dimension is [8, Theorems 2.7.2,Theorem 4.5.1], and the general case of PEL type AC is proven in [54, Theorems 4.3.10,4.3.12]. ∎
@@
[8, 54] The partially minimally compactified Igusa variety $\mathrm{Ig}^{b,*}$ is affine; in particular, the attached adic space $\mathfrak{Ig}^{b,*}_{C}$ is affinoid perfectoid. Moreover, there exists a proper map
@@
For the affineness, the case of PEL type $A$ attached to a global quasi-split unitary group of even dimension is covered by [8, Theorem 1.7] and [8, Lemma 4.5.2]. The general case of PEL type AC is covered in [54, Lemma 3.3.7]. To see the properness, we note that the map
@@
is the one appearing in the Stein factorization described in [8, Proposition 3.3.4] and [54, Proposition 3.3.5]. This also shows that one has an isomorphism on global sections. ∎
@@
The fact that $\mathfrak{Ig}_{C}^{b,*}$ is affinoid perfectoid follows from the previous Theorem. Moreover, $\partial\mathfrak{Ig}_{C}^{b,*}$ is a Zariski closed subspace, since it came from considering the adic generic fiber of a formal model of the perfect closed subscheme $\partial\mathrm{Ig}^{b,*}\subset\mathrm{Ig}^{b,*}$, and a Zariski closed subspace of a perfectoid is still perfectoid by [3, Remark 7.5]. The claim for the toroidal compactification follows since the map $\mathfrak{Ig}_{C}^{b,\mathrm{tor}}\rightarrow\mathfrak{Ig}_{C}^{b,*}$ is proper, and maps the boundary $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}$ to $\partial\mathfrak{Ig}_{C}^{b,*}$. ∎
@@
For $x:\mathop{\rm Spa}(C,O_{C})\rightarrow\mathcal{F}\ell_{G,\mu^{-1}}^{b,\Diamond}$ a geometric point, we have isomorphisms
@@
induced by the natural open immersions $\pi_{\mathrm{HT}}^{-1}(x)\hookrightarrow(\pi_{\mathrm{HT}}^{*})^{-1}(x)\simeq\overline{\mathfrak{Ig}}_{C}^{b,*}$ (resp. $\pi_{\mathrm{HT}}^{-1}(x)\hookrightarrow(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)\simeq\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$), as given by Theorem 3.1. Here $\overline{\partial\mathfrak{Ig}}_{C}^{b,*}$ (resp. $\overline{\partial\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$) is the Zarisiki closed subset of $\overline{\mathfrak{Ig}}_{C}^{b,*}$ (resp. $\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$) defined by the canonical compactification of the boundary $\partial\mathfrak{Ig}^{b}_{C}\subset\mathfrak{Ig}^{b}_{C}$ (resp. $\partial\mathfrak{Ig}_{C}^{b,\mathrm{tor}}\subset\mathfrak{Ig}^{b,\mathrm{tor}}_{C}$).
@@
obtained by base-changing the closed immersion $\partial\mathcal{S}^{\mathrm{tor}}_{K^{p},C}\hookrightarrow\mathcal{S}^{\mathrm{tor}}_{K^{p},C}$ to the fiber $(\pi_{\mathrm{HT}}^{\mathrm{tor}})^{-1}(x)$ and applying Theorem 3.1. To show the desired claim, it suffices to show this is an isomorphism. We note that both the RHS and LHS are perfectoid by [3, Remark 7.5], since they are Zariski closed in the perfectoid space $\overline{\mathfrak{Ig}}_{C}^{b,\mathrm{tor}}$. Moreover, both the LHS and RHS are partially proper; therefore, to show this map is an isomorphism, it suffices to show it induces an isomorphism on rank one geometric points using [56, Lemma 5.4] (where we note that condition (iv) of [56, Lemma 5.4] will be implied by the claim on rank one points and the partial properness). In particular, given $C^{\prime}/C$ a complete algebraically closed non-archimedean field, we claim that there exists a Cartesian diagram of the form